OpenStudy (anonymous):

I need help on this problem I don't understand how to factor this a^2-6a+8

OpenStudy (anonymous):

is it factoring?

OpenStudy (anonymous):

(a-4) (a-2)

OpenStudy (anonymous):

yes its factoring I see the a^2 and the 8 but where is the 6a?

OpenStudy (anonymous):

Carolinejean are you still there?

OpenStudy (anonymous):

dude thats the answer.

OpenStudy (anonymous):

think of if this way ... the format of a quadratic is \[ax^2+bx+c=0\] right ? to factor, think of what 2 numbers multiply to make up c or in this case +8, and at the same time add to b or in this case -6 ... \[-2*-4=8\] and \[-2 + -4 =-6\]... so we write \[(x-2)(x-3)=0\] and to solve, since it is a multiplication of 2 terms, then we should have 2 cases where this is valid :\[(x-2)=0 ==> x=2\] or \[(x-3)=0 ==> x=3\] voila!